Convert the following to polar form, answer in radians
(i) 1 + j3
(ii)5 - j2

Answer by Edwin McCravy(20086)

Convert the following to polar form, answer in radians
(i) 1 + j3
(ii)5 - j2

In general, the picture of a + jb is the vector that 
connects the origin (0,0) to the point (a,b)  

Therefore, the picture of 1 + j3 is the vector that 
connects the origin to the point (1,3) is this vector:



Now we draw in the y-coordinate:



That triangle has base 1, and the green side is 3, so the vector is 
the hypotenuse and that equals 
 _____    ___    __
Ö1²+3² = Ö1+9 = Ö10

The angle between the vector and the x-axis has a tangent of 3/1, so
the angle is 1.249 radians, approximately.  So the polar form can be 
written any of the following ways:
 __
Ö10[sin(1.249) + j·cos(1.249)]
 __
Ö10 cjs(1.249)
 __
Ö10 Ð 1.249
 
------   

(ii)5 - j2

The picture of 5 - j2 is the vector that 
connects the origin to the point (5,-2) is this vector:



Now we draw in the y-coordinate:



That triangle has top side 5, and the green side is -2, so the vector 
is the hypotenuse and that equals 
 ________    ____    __
Ö5²+(-2)² = Ö25+4 = Ö29

The angle between the vector and the x-axis has a tangent of -2/5, so
the angle is -.381 radians, approximately.  So the polar form can be 
written any of the following ways:
 __
Ö29[sin(-.381) + j·cos(-.381)]
 __
Ö29 cjs(-.381)
 __
Ö29 Ð -.381

Edwin